# Heralded Optical Entanglement Generation via the Graph Picture of Linear Quantum Networks

### v2

@@ V Discussions
We can extend our discussion to qudit entanglements such as qudit $N$-partite GHZ states that we can also generate with the sculpting protocol as given in Ref. [16]. We can choose another photonic degree of freedom as the internal state for encoding qudits, e.g., the orbital angular momentum (OAMs). For the case, the PBSs are replaced with OAM beam splitters [31] and HWPs with OAM-only Fourier transformation operators [34].
Since all our schemes use feasible single-photon sources and linear optical operators, we anticipate that they will be validated through real optical experiments in the near future. Alternatively, one can also use online quantum computing platforms such as Quandella [29] that offer the linear operations involving ideal photons by adopting our circuits into dual-rail encoding circuits.
@@ Acknowledgements
SC is grateful to Prof. Ana Belen Sainz and Prof. Jung-Hoon Chun for their support on this research. This research is funded by National Research Foundation of Korea (NRF, RS-2023-00245747, 2021M3H3A103657313, 2022M3K4A1094774, and 2023M3K5A1094805), Korea Institute of Science and Technology (2E31021), and Foundation for Polish Science (IRAP project, ICTQT, contract no.2018/MAB/5, co-financed by EU within Smart Growth Operational Programme)
Appendix A GHZ state generation scheme
@@
    $\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,H}{\hat{a}}^{\dagger}_{j,V}~{}\to~{}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}$ (39)
@@
    $\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}~{}\to~{}\frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})$ (40)
@@
    $\frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})~{}\to~{}\frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{(j\oplus_{N}1)2,V})$ (41)
@@
    $\frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{(j\oplus_{N}1)2,V})~{}\to~{}$
@@
The success probability is $\frac{1}{2^{2N}}$ without feed-forward, which becomes $\frac{1}{2^{2N-1}}$ with feed-forward.
Appendix B W state generation scheme
@@
    $\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,H}{\hat{a}}^{\dagger}_{j,V}\Big{)}{\hat{a}}^{\dagger}_{X,H}~{}\to~{}\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}$ (46)
@@
    $\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}~{}\to~{}\frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{Xk,D}$ (47)
@@
    $\frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{Xk,D}~{}\to~{}\frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{Xj,V})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{k3,D}$ (48)
@@
    $\to~{}\frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,D}-{\hat{a}}^{\dagger 2}_{Xj,A})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{k3,V}$
@@
Appendix C $N=3$ Type 5 entangled state generation scheme
@@
    $\Big{(}\prod^{3}_{j=1}{\hat{a}}^{\dagger}_{j,H}{\hat{a}}^{\dagger}_{j,V}\Big{)}{\hat{a}}^{\dagger}_{X,H}{\hat{a}}^{\dagger}_{Y,H}{\hat{a}}^{\dagger}_{Z,H}~{}\to~{}\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}{\hat{a}}^{\dagger}_{Y,D}{\hat{a}}^{\dagger}_{Z,D}$ (53)
@@
    $\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}{\hat{a}}^{\dagger}_{Y,D}{\hat{a}}^{\dagger}_{Z,D}~{}\to~{}\frac{1}{24\sqrt{2}}\prod^{3}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})({\hat{a}}^{\dagger}_{X1,D}+{\hat{a}}^{\dagger}_{X2,D})\sum_{k=1}^{3}{\hat{a}}^{\dagger}_{Yk,D}\sum_{l=1}^{3}{\hat{a}}^{\dagger}_{Zl,D}$ (54)
@@
    ×$~{}~{}~{}~{}\times({\hat{a}}^{\dagger}_{13,D}+{\hat{a}}^{\dagger}_{Y2,D})({\hat{a}}^{\dagger}_{23,D}+{\hat{a}}^{\dagger}_{Y3,D}+{\hat{a}}^{\dagger}_{Z2,D})({\hat{a}}^{\dagger}_{33,D}+{\hat{a}}^{\dagger}_{X2,D}+{\hat{a}}^{\dagger}_{Z3,D})$ (55)
@@
    ×$~{}~{}~{}~{}\times({\hat{a}}^{\dagger}_{13,V}+{\hat{a}}^{\dagger}_{Y2,H})({\hat{a}}^{\dagger}_{23,V}+{\hat{a}}^{\dagger}_{Y3,H}+{\hat{a}}^{\dagger}_{Z2,H})({\hat{a}}^{\dagger}_{33,V}+{\hat{a}}^{\dagger}_{X2,H}+{\hat{a}}^{\dagger}_{Z3,H})$
@@
    ×$~{}~{}~{}~{}\times\Big{(}({\hat{a}}^{\dagger}_{31,H}+{\hat{a}}^{\dagger}_{32,V})^{2}-({\hat{a}}^{\dagger 2}_{Z1,H}-{\hat{a}}^{\dagger 2}_{3,V})^{2}\Big{)}({\hat{a}}^{\dagger}_{13,V}+{\hat{a}}^{\dagger}_{Y2,H})({\hat{a}}^{\dagger}_{23,V}+{\hat{a}}^{\dagger}_{Y3,H}+{\hat{a}}^{\dagger}_{Z2,H})({\hat{a}}^{\dagger}_{33,V}+{\hat{a}}^{\dagger}_{X2,H}+{\hat{a}}^{\dagger}_{Z3,H})$ (56)
@@
    $~{}~{}~{}~{}~{}~{}+{\hat{a}}^{\dagger}_{X1,H}{\hat{a}}^{\dagger}_{1,V}{\hat{a}}^{\dagger}_{21,H}{\hat{a}}^{\dagger}_{22,V}{\hat{a}}^{\dagger}_{Z1,H}{\hat{a}}^{\dagger}_{3,V}{\hat{a}}^{\dagger}_{13,V}{\hat{a}}^{\dagger}_{Y3,H}{\hat{a}}^{\dagger}_{33,V}+{\hat{a}}^{\dagger}_{X1,H}{\hat{a}}^{\dagger}_{1,V}{\hat{a}}^{\dagger}_{Y1,H}{\hat{a}}^{\dagger}_{2,V}{\hat{a}}^{\dagger}_{31,H}{\hat{a}}^{\dagger}_{32,V}{\hat{a}}^{\dagger}_{13,V}{\hat{a}}^{\dagger}_{23,V}{\hat{a}}^{\dagger}_{Z3,H}$
    $~{}~{}~{}~{}~{}+{\hat{a}}^{\dagger}_{X1,H}{\hat{a}}^{\dagger}_{1,V}{\hat{a}}^{\dagger}_{Y1,H}{\hat{a}}^{\dagger}_{2,V}{\hat{a}}^{\dagger}_{Z1,H}{\hat{a}}^{\dagger}_{3,V}{\hat{a}}^{\dagger}_{13,V}{\hat{a}}^{\dagger}_{23,V}{\hat{a}}^{\dagger}_{33,V}\Big{)}$ (57)
@@
Appendix D From the polarization encoding to the dual-rail encoding

### v3

@@ 5 Discussions
We can extend our discussion to qudit entanglements such as qudit $N$-partite GHZ states [36] that we can also generate with the sculpting protocol as given in Ref. [16]. We can choose another photonic degree of freedom as the internal state for encoding qudits, e.g., the orbital angular momentum (OAMs). For the case, the PBSs are replaced with OAM beam splitters [31] and HWPs with OAM-only Fourier transformation operators [37].
@@ Acknowledgements
SC is grateful to Prof. William J. Munro, Prof. Ana Belen Sainz, and Prof. Jung-Hoon Chun for their support on this research. This research is funded by National Research Foundation of Korea (NRF, RS-2023-00245747, 2021M3H3A103657313, 2022M3K4A1094774, and 2023M3K5A1094805), Korea Institute of Science and Technology (2E31021), and Foundation for Polish Science (IRAP project, ICTQT, contract no.2018/MAB/5, co-financed by EU within Smart Growth Operational Programme)
Appendix A Undirected bipartite graph representation of bosonic systems in LQG picture
As we can see from Section 2.1, essential information on generating multipartite entanglement with sculpting protocol lies in the sculpting operators. Therefore, we can always consider a simpler graph representation that only reflects the property of sculpting operators without the direction of edges, i.e., undirected bipartite representation [16]. Here we briefly summarize the undirected representation of LQG picture and how they are directly related to the directed representation.
Boson systems with sculpting operator Bipartite Graph $G_{b}=(U\cup V,E)$
Spatial modes Labelled vertices $\in$ $U$
$\hat{A}^{(l)}$ ($l\in\{1,2,\cdots,N\}$) Unlabelled vertices $\in$ $V$
Spatial distributions of $\hat{A}^{(l)}$ Edges $\in$ $E$
Probability amplitude $\alpha_{j}^{(l)}$ Edge weight $\alpha_{j}^{(l)}$
Internal state $\psi_{j}^{(l)}$ Edge weight $\psi_{j}^{(l)}$
Table 2: Correspondence relations from a sculpting operator to a undirected bipartite graph
Then the identities
${\hat{a}}_{j,+}{\hat{a}}_{j,-}{\hat{a}}^{\dagger}_{j,0}{\hat{a}}^{\dagger}_{j,1}|vac\rangle=0,$
${\hat{a}}_{j,0}^{n}{\hat{a}}^{\dagger}_{j,0}{\hat{a}}^{\dagger}_{j,1}|vac\rangle={\hat{a}}_{j,1}^{n}{\hat{a}}^{\dagger}_{j,0}{\hat{a}}^{\dagger}_{j,1}|vac\rangle=0.\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (n\geq 2)$ (38)
are translated into the undirected bigraph language as
[Uncaptioned image].[Uncaptioned $\end{gathered}.$ (40)
Note that the above graphical relations correspond to (10) and (12) of the directed bigraph representation.
Undirected EPM bigraphs are now defined as follows. If all the edges of a bigraph attach to the circles as one of the following undirected bigraphs
[Uncaptioned image],[Uncaptioned $\end{gathered},$ (42)
then it is an undirected EPM bigraph.
Three undirected EPM biraphs that corresponds to (a) GHZ state, (b) W-state, (c) $N=3$ Type 5 entangled state are proposed in Ref. [16] as follows:
[Uncaptioned image][Uncaptioned $\end{gathered}$ (44)
Appendix B GHZ state generation scheme
@@
    $\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,H}{\hat{a}}^{\dagger}_{j,V}\leavevmode\nobreak\ \to\leavevmode\nobreak\ \prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}$ (46)
@@
    $\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\leavevmode\nobreak\ \to\leavevmode\nobreak\ \frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})$ (47)
@@
    $\frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})\leavevmode\nobreak\ \to\leavevmode\nobreak\ \frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{(j\oplus_{N}1)2,V})$ (48)
@@
    $\frac{1}{2^{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{(j\oplus_{N}1)2,V})\leavevmode\nobreak\ \to\leavevmode\nobreak$
@@
The success probability is determined by the normalization factor of the final state (50) after the postselection. The success probability is $\frac{1}{2^{2N}}$ without feed-forward, which becomes $\frac{1}{2^{2N-1}}$ with feed-forward.
Appendix C W state generation scheme
@@
    $\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,H}{\hat{a}}^{\dagger}_{j,V}\Big{)}{\hat{a}}^{\dagger}_{X,H}\leavevmode\nobreak\ \to\leavevmode\nobreak\ \Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}$ (53)
@@
    $\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}\leavevmode\nobreak\ \to\leavevmode\nobreak\ \frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{Xk,D}$ (54)
@@
    $\frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{Xk,D}\leavevmode\nobreak\ \to\leavevmode\nobreak\ \frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{Xj,V})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{k3,D}$ (55)
@@
    $\to\leavevmode\nobreak\ \frac{1}{2^{N}\sqrt{N}}\prod^{N}_{j=1}({\hat{a}}^{\dagger 2}_{j1,D}-{\hat{a}}^{\dagger 2}_{Xj,A})\sum_{k=1}^{N}{\hat{a}}^{\dagger}_{k3,V}$
@@
Appendix D $N=3$ Type 5 entangled state generation scheme
@@
    $\Big{(}\prod^{3}_{j=1}{\hat{a}}^{\dagger}_{j,H}{\hat{a}}^{\dagger}_{j,V}\Big{)}{\hat{a}}^{\dagger}_{X,H}{\hat{a}}^{\dagger}_{Y,H}{\hat{a}}^{\dagger}_{Z,H}\leavevmode\nobreak\ \to\leavevmode\nobreak\ \Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}{\hat{a}}^{\dagger}_{Y,D}{\hat{a}}^{\dagger}_{Z,D}$ (60)
@@
    $\Big{(}\prod^{N}_{j=1}{\hat{a}}^{\dagger}_{j,D}{\hat{a}}^{\dagger}_{j,A}\Big{)}{\hat{a}}^{\dagger}_{X,D}{\hat{a}}^{\dagger}_{Y,D}{\hat{a}}^{\dagger}_{Z,D}\leavevmode\nobreak\ \to\leavevmode\nobreak\ \frac{1}{24\sqrt{2}}\prod^{3}_{j=1}({\hat{a}}^{\dagger 2}_{j1,H}-{\hat{a}}^{\dagger 2}_{j2,V})({\hat{a}}^{\dagger}_{X1,D}+{\hat{a}}^{\dagger}_{X2,D})\sum_{k=1}^{3}{\hat{a}}^{\dagger}_{Yk,D}\sum_{l=1}^{3}{\hat{a}}^{\dagger}_{Zl,D}$ (61)
@@
    ×$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \times({\hat{a}}^{\dagger}_{13,D}+{\hat{a}}^{\dagger}_{Y2,D})({\hat{a}}^{\dagger}_{23,D}+{\hat{a}}^{\dagger}_{Y3,D}+{\hat{a}}^{\dagger}_{Z2,D})({\hat{a}}^{\dagger}_{33,D}+{\hat{a}}^{\dagger}_{X2,D}+{\hat{a}}^{\dagger}_{Z3,D})$ (62)
@@
    ×$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \times({\hat{a}}^{\dagger}_{13,V}+{\hat{a}}^{\dagger}_{Y2,H})({\hat{a}}^{\dagger}_{23,V}+{\hat{a}}^{\dagger}_{Y3,H}+{\hat{a}}^{\dagger}_{Z2,H})({\hat{a}}^{\dagger}_{33,V}+{\hat{a}}^{\dagger}_{X2,H}+{\hat{a}}^{\dagger}_{Z3,H})$
@@
    ×$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \times\Big{(}({\hat{a}}^{\dagger}_{31,H}+{\hat{a}}^{\dagger}_{32,V})^{2}-({\hat{a}}^{\dagger 2}_{Z1,H}-{\hat{a}}^{\dagger 2}_{3,V})^{2}\Big{)}({\hat{a}}^{\dagger}_{13,V}+{\hat{a}}^{\dagger}_{Y2,H})({\hat{a}}^{\dagger}_{23,V}+{\hat{a}}^{\dagger}_{Y3,H}+{\hat{a}}^{\dagger}_{Z2,H})({\hat{a}}^{\dagger}_{33,V}+{\hat{a}}^{\dagger}_{X2,H}+{\hat{a}}^{\dagger}_{Z3,H})$ (63)
@@
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +{\hat{a}}^{\dagger}_{X1,H}{\hat{a}}^{\dagger}_{1,V}{\hat{a}}^{\dagger}_{21,H}{\hat{a}}^{\dagger}_{22,V}{\hat{a}}^{\dagger}_{Z1,H}{\hat{a}}^{\dagger}_{3,V}{\hat{a}}^{\dagger}_{13,V}{\hat{a}}^{\dagger}_{Y3,H}{\hat{a}}^{\dagger}_{33,V}+{\hat{a}}^{\dagger}_{X1,H}{\hat{a}}^{\dagger}_{1,V}{\hat{a}}^{\dagger}_{Y1,H}{\hat{a}}^{\dagger}_{2,V}{\hat{a}}^{\dagger}_{31,H}{\hat{a}}^{\dagger}_{32,V}{\hat{a}}^{\dagger}_{13,V}{\hat{a}}^{\dagger}_{23,V}{\hat{a}}^{\dagger}_{Z3,H}$
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +{\hat{a}}^{\dagger}_{X1,H}{\hat{a}}^{\dagger}_{1,V}{\hat{a}}^{\dagger}_{Y1,H}{\hat{a}}^{\dagger}_{2,V}{\hat{a}}^{\dagger}_{Z1,H}{\hat{a}}^{\dagger}_{3,V}{\hat{a}}^{\dagger}_{13,V}{\hat{a}}^{\dagger}_{23,V}{\hat{a}}^{\dagger}_{33,V}\Big{)}$ (64)
@@
Appendix E From the polarization encoding to the dual-rail encoding